From GlueXWiki
Jump to: navigation, search

April 11, 2013 FDC meeting


  1. Production Construction Tracking (Dave)
    • Package #3 repair
  2. Installation (Lubomir, Bill)
  3. Engineering (Bill)
    • Cooling system tests
  4. Electronics (Chris, Nick)
  5. Tests at EEL126 FDC E-log (Beni)
  6. Other


Participants: Eugene, Fernando, Dave, Chris, Nick, Simon, Beni, and Lubomir.


- Dave for the spare package: Two wire planes (#2 and #3) have been deadened, two more remaining. The filed wire on #2 was fixed (broke during deadening). It turned out we don't have enough HV cables (short cables with male connectors) that are soldered on the wire frames; we have to order them. This is because we originally planned for 4 spare cells.

- Package #3 repair: the three field wires on the wire frame of cell #6 were fixed and on Monday we installed the sixth cell on the package upside-down. We ended up holding the spacer ring with fingers (taping to polyethylene didn't work) when installing it together with the wire frame, but it was successful: no gas leakage was found. Since Monday we are flushing the package with gas, oxygen went down to 700ppm and today we applied 2000V/-500V so far without problems. We will continue testing it till Monday, but already tomorrow we can start replacing the cooling tubes - they have to be positioned in a different way as for the other packages. Next week we plan to finish also with the grounding connections, HV connections and with the installation of the manifolds. Then we will move it (during the week of 22-26 of April) to 126 for testing with the full electronics. Fernando: by that time we will have 30 fADC125 ready to be used (3 will go to CMU).


- The channel assignment schemes are attached above. Fernando explained how to use the first one, the channels/cables are ordered according the their rack, crate, slot numbers. The other two files (Vlad's) represent the view from the detector side, separately for the signal and LV cables. The cables are ordered in increasing azimuthal angle, started from the top counterclockwise when looking from the upstream side.

- The schedule for the FDC installation and related activities (linked above) was updated from the recent version 3 of the rebaselined schedule. There are important changes related to the electronics: the installation of the fADC125s in the racks is postponed by ~8 weeks, starts Nov 23. If this is the case, the FDC installation has to be postponed as well. Eugene: this was done because of the manpower spike for this period of time. According to Fernando there are no problems with the old schedule and he has the people to do the job, so Eugene will discuss this with the management.

- Practicing the cable installation: the mesh cylinder (one half) will be installed on one of the installation carts at Blue Crab; Casey got the parts for that. On another cart we will install one of the packages and the mock-up to simulate the three other packages; for the mock-up we need to find brackets to attach it to the rails. Chris and Nick already brought 132 signal cables (a quoter of the total number) to Blue Crab.


- Casey finished the set-up for the one-loop test with the SMC chiller at 126. It turned out the water chiller outputs 260psi and the SMC chiller requires <100psi, so Bill has to make modifications before we can do the test. For the package test at Blue Crab Bill made modifications to avoid the leakage in the testing set-up. There we use CO2 at 100psi, and both Fernando and Eugene were concerned about the safety; will be discussed with Bill.

- There was an idea to monitor the positions of the packages when inside the magnet, using telescopes attached to the face of the magnet looking at the optical targets. Eugene: the z-position is much more important than the x-y, therefore he thinks it's not worth doing it now; there are also no money for that in the schedule.


- Fernando once more explained the quality issues with the fADC125. Since some of the modules were already fixed by the company he expects to have 30 modules available very soon. For the full electronics test, Fernando wants to see if everything is OK with the cooling before powering the pre-amps.

- Nick will give us 22 pre-amps for the cooling system tests. They will be powered from the LV panel at 126.

Tests at EEL126

- Package #2 was moved yesterday from 126 to Blue Crab.

- Beni: when testing package #2 with lower flow (100ccpm instead of the usual 200ccpm), some of the channels were tripping once per day. It could be some kind of gas separation, or very low flow in some of the cells due to different impedances (Beni), or some impurities in the gas/chamber (Fernando), or something else?